A population consists of the five measurements

 2 , 6 , 8 , 0 , and 1 . How many different samples 

 of size n = 2 can be drawn from the population.

 Construct a table for the possible samples and

 sample mean.        

  



1
Expert's answer
2021-06-08T12:12:23-0400

1.

We have population values "2,6,8,0,1" population size "N=5" and sample size "n=2."Thus, the number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{5}{2}=10"


2.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 0,1 & 0.5 \\
 \hdashline
 2 & 0,2 & 1 \\
 \hdashline
 3 & 0,6 & 3 \\
 \hdashline
 4 & 0,8 & 4\\
 \hdashline
 5 & 1,2 & 1.5 \\
\hdashline
 6 & 1,6 & 3.5 \\
 \hline
7 & 1,8 & 4.5 \\
 \hline
 8 & 2,6 & 4 \\
 \hline
9 & 2,8 & 5 \\
 \hline
10 & 6,8 & 7 \\
 \hline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 0.5 & 1 & 0.1 & 0.05 & 0.025 \\
 \hdashline
 1.0 & 1 & 0.1 & 0.10 & 0.1 \\
 \hdashline
 1.5 & 1 & 0.1 & 0.15 & 0.225 \\
 \hdashline
 3.0 & 1 & 0.1 & 0.3 & 0.9 \\
 \hdashline
 3.5 & 1 & 0.1 & 0.35 & 1.225 \\
 \hdashline
 4.0 & 2& 0.2 & 0.80 & 3.2\\
 \hdashline
 4.5 & 1& 0.1 & 0.45 & 2.025 \\
 \hdashline
5.0 & 1& 0.1 & 0.5 & 2.5 \\
 \hdashline
7.0 & 1& 0.1 & 0.7 & 4.9 \\
 \hdashline
 Total & 10 & 1 & 3.4 & 15.1 \\ \hline
\end{array}"



3.

Mean


"\mu=\dfrac{2+6+8+0+1}{5}=3.4"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=3.4"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=3.4=\mu"


4.

Variance


"\sigma^2=\dfrac{1}{5}\big((2-3.4)^2+(6-3.4)^2+(8-3.4)^2"




"+(0-3.4)^2+(1-3.4)^2\big)=9.44"


Standard deviation



"\sigma=\sqrt{\sigma^2}=\sqrt{9.44}\approx3.072458"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=15.1-(3.4)^2=3.54"




"\sqrt{Var(\bar{X})}=\sqrt{3.54}\approx1.881489"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{9.44}{2}(\dfrac{5-2}{5-1})"




"=3.54, True"